OpenStudy (anonymous):

a hydrogen atom has a diameter of about 10 nm express this diameter in micrometers, meters, and milimeters

OpenStudy (goformit100):

Ya for that you will have to code something. It is as follows :- This is what I was talking about... first you will have to Memorize this table http://nobel.scas.bcit.ca/chem0010/tutorials/SI%20units,%20prefixes.pdf I too have memorized it When I was in My 12th grade.

OpenStudy (abb0t):

Well nanometers is 1 nanometer = 1 \(\times\)10\(^{-9}\)m 1 micrometer = 1 \(\times\)10\(^{-6}\) m 1 millimeter = 1 \(\times\)10\(^{-3}\)m